OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

OpenStudy (mr.math):

LOL!

OpenStudy (anonymous):

I'm trying to help.

OpenStudy (anonymous):

Well i'm stupid and dont get it so no worth trying

OpenStudy (anonymous):

Let's take it one step at a time. First, re-write that equation with '6' instead of 'm'. What do you get?

OpenStudy (anonymous):

I don't think you're stupid, you just are having a hard time with math.

OpenStudy (anonymous):

The general form, remember, is: y-y0 = m(x-x0). Change 'm' to '6', what do you get?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

do you see the place in that equation where there is an 'm'?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So change it to a 6. Now what does that equation look like?

hero (hero):

ixl.com

OpenStudy (anonymous):

honestly.. i dont know any of it. im failing algerbra.

OpenStudy (anonymous):

Erika just re-type the equation, but with '6' instead of 'm'. That's all I'm asking you to do.

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

just re-type the equation, but with '6' instead of 'm'. That's all I'm asking you to do

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. where do i put the m

OpenStudy (anonymous):

y-y0 = m(x-x0). Just change 'm' to '6', what do you get?

OpenStudy (anonymous):

'm' is what stands for 'slope'. That's why when they tell you the slope is 6 you know that's what you put for 'm'.

OpenStudy (anonymous):

okay so the 6 goes where the m is?

OpenStudy (anonymous):

right. what do you get?